if a body is at rest. what is the ratio of total distance travelled in 3rd and 4th sec????????

Grade:10

1 Answers

Vikas TU
7 years ago
From the eqn. s=u+a/2(2n-1)
 
for 3rd second,
s3 = 0 + (a/2)*(5) = > 5a/2
 
for 4th second,
s4 = 0 + (a/2)*(7) = > 7a/2
 
Now,
s3/s4 =  5a/7a = 5/7 => 5:7